OpenStudy (anonymous):

A pair of dice is rolled. What is the probability of getting a sum greater than 3? and a sum less than 10?

OpenStudy (apoorvk):

well, a sum less that 3 can be 2. And the no. of ways to get the sum as 2 is only one --> (1,1) so, one favorable outcome. sample space = 36. so what's the probability?

OpenStudy (anonymous):

(2,1) ?

OpenStudy (anonymous):

or (1,2) ?

OpenStudy (anonymous):

so is it 35/36? since 1 outcome is less then 3

OpenStudy (apoorvk):

Oh am sorry, we need a sum *greater* than 3. Misread it. :/ Yes, you were going right, so we include (1,2) and (2,1) with(1,10 in the set of "unfavorable outcomes". so, favorable outcomes = 36 - 3 = 33 hence probability = 33/36

OpenStudy (apoorvk):

As for the second part, for a sum 'lesser' than 10, we use the same method to find out the no. of favorable outcomes. Outcomes with sum equal to or greater than 10 = (4,6), (6,4), (5,5), (5,6), (6,5) and (6,6). = 6 so no of outcomes with sum less than 10 = 36 - 6 = 30 Hence probability = 30/36 = 5/6

OpenStudy (apoorvk):

@monroe17